OpenStudy (anonymous):

Solve. 1/6r+1/4=1/3+3/4r

OpenStudy (mathstudent55):



OpenStudy (zehanz):

Your equation:\[\frac{ 1 }{ 6 }r+\frac{ 1 }{ 4 }=\frac{ 1 }{ 3 }+\frac{ 3 }{ 4 }r\] On the other hand, you could have:\[\frac{ 1 }{ 6r }+\frac{ 1 }{ 4 }=\frac{ 1 }{ 3 }+\frac{ 3 }{ 4r }\] Apart from this, in both cases the same method works: try to find a number to multiply everything with, so that the fractions are gone.

OpenStudy (zehanz):

Suppose it is the first case, look at the numerators: 12 can be divided by everyone of them, so multiply verything with 12 and see what you get...

OpenStudy (anonymous):

yes @mathstudent55

OpenStudy (zehanz):

Then you try my hint ;)

OpenStudy (anonymous):

I really don't understand this..

OpenStudy (mathstudent55):

Look at all the denomionators: 6, 4, 3, 4 What is the least common multiple (LCM) of the denominators? It's 12 Multiply both sides by the LCM